# Evaluating Large Language Model based Personal Information Extraction and Countermeasures

### v1

@@ 5.2 Real-world Datasets
As complementary to our synthetic dataset, we collected two real-world datasets to assess the performance of LLM-based PIE in a more practical and realistic scenario. To mitigate the ethical and privacy concerns, we collect personal profiles that are already in the public domain. The first real-world dataset is collected from the "List of Top 100 Famous People" website [11], which contains personal profiles of famous people from the nineteenth, twentieth, or twenty-first centuries. The list includes famous actors, politicians, entrepreneurs, writers, artists, and humanitarians. The second dataset is derived from "Wikipedia’s 19th century American physicians" [10], which contains personal profiles of known American physicians in the 19th century. All personal profiles in both datasets are in HTML format. In addition, we manually collect the ground-truth personal information in these two real-world datasets. We note that some personal profiles may not have some categories of personal information.
The use of these two real-world datasets is beneficial to our study. In particular, it enables us to evaluate LLM-based PIE in realistic scenarios where the data has a more complicated structure and thus less predictable than the synthetic dataset. This could help understand the applicability of LLMs in extracting personal information in reality.
@@
API-based PaLM 2 text-bison-001 540B Google
@@
6 Evaluating Attacks
6.1 Experimental Settings
6.1.1 LLMs
We leverage the following LLMs to evaluate the attacks: PaLM 2 text-bison-001 [15], PaLM 2 chat-bison-001 [15], gemini-pro [46], Flan-UL2 [50, 45, 44], Vicuna-13b-v1.3 [21, 54], Vicuna-7b-v1.3 [21], OpenAI’s gpt-3.5-turbo [6], gpt-4 [34], Llama-2-7b-chat [48, 8], and InternLM-Chat-7B [47]. Table 2 shows the details of those LLMs. Unless otherwise mentioned, we use the PaLM 2 text-bison-001 as the default model, due to its free-of-charge APIs and powerful capabilities (with 540B parameters), enabling us to conduct systematic evaluation.
6.1.2 Attack settings
We use the following settings for two components of the attack framework:
    Table 3: The effectiveness of LLMs at extracting 8 categories of personal information on our three datasets. The results are presented as Accuracy / BERT score for email addr. and Tel., and as Rouge-1 score / BERT score for other categories.
@@ 6.1.2 Attack settings
    vicuna-13b-v1.3 Synthetic 62% / 0.62 56% / 0.56 66% / 0.65 100% / 1.00 62% / 0.44 50% / 0.39 65% / 0.66 69% / 0.66
@@
    vicuna-7b-v1.3 Synthetic 55% / 0.55 89% / 0.89 49% / 0.46 98% / 0.98 56% / 0.38 49% / 0.37 85% / 0.83 75% / 0.68
@@
    llama-2-7b-chat-hf Synthetic 31% / 0.31 50% / 0.50 27% / 0.26 34% / 0.34 39% / 0.24 42% / 0.29 38% / 0.36 17% / 0.14
@@
    internlm-chat-7b Synthetic 96% / 0.96 96% / 0.96 83% / 0.82 100% / 1.00 55% / 0.37 51% / 0.40 64% / 0.67 60% / 0.51
@@
    Regular expression 100% / 1.00 36% / 0.36 49% / 0.41 96% / 0.96 20% / 0.12 15% / 0.06 5% / 0.12 1% / 0.05
    Keyword search 13% / 0.13 17% / 0.17 10% / 0.10 3% / 0.03 42% / 0.34 39% / 0.29 5% / 0.12 1% / 0.05
    Entity detection 1% / 0.01 24% / 0.24 17% / 0.10 8% / 0.08 0.22% / 0.00 0.22% / 0.00 4% / 0.11 1% / 0.05
    ML method 77% / 0.77 8% / 0.08 10% / 0.10 39% / 0.39 0.22% / 0.00 0.22% / 0.00 5% / 0.12 1% / 0.05
    Table 5: Impact of prompt style on the effectiveness of LLM-based PIE.
    Prompt Type Personal Information
@@
    Direct 93% / 0.93 95% / 0.95 77% / 0.67 98% / 0.96 67% / 0.53 76% / 0.59 80% / 0.70 51% / 0.37
    Persona 98% / 0.98 78% / 0.78 73% / 0.68 100% / 1.00 56% / 0.42 72% / 0.53 80% / 0.73 51% / 0.38
    Contextual 93% / 0.93 95% / 0.95 75% / 0.54 97% / 0.96 62% / 0.47 74% / 0.56 83% / 0.74 50% / 0.35
    Pseudocode 100% / 1.00 99% / 0.99 81% / 0.74 100% / 1.00 52% / 0.33 71% / 0.53 82% / 0.75 51% / 0.36
  • Personal profile processing:  We conduct the experiments on the three datasets described in Section 5. Unless otherwise mentioned, we use the Synthetic dataset as the default one. Given a personal profile, an attacker aims to extract the following categories of personal information: name, email address, phone number, mailing address, work experience, educational experience, affiliation, and occupation. By default, we leverage an HTML parser to remove redundant information (see Section 4.3 for details).
@@ 6.1.3 Compared baselines
We adapt several traditional PIE methods as baselines, including regular expression, keyword Search, entity detection, and ML-based method. Their implementation details are as follows:
  • Regular expression: For each category of personal information, we develop corresponding regular expressions based on the possible structures of the corresponding texts. We use these regular expressions to match and extract the personal information from a personal profile. Take email address as an example. Following the typical format of an email, the regular expression can be represented as “[a-zA-Z0-9._%+-]+@[a-zA-Z0-9.-]+\.[a-zA-Z]{2,}.”. Table 14 in the Appendix shows the specific regular expression for each category of personal information.
  • Keyword search: We first search for the keywords (e.g., “email”) present in the HTML titles. Then, we extract the text located under the identified HTML titles by keyword search. Finally, we post-process the extracted text by removing the redundant or unrelated information.
  • Entity detection: We utilize spaCy [1], a popular tool used for named entity recognition, to extract personal information from a personal profile. The implementation of spaCy has achieved notable popularity, with a noteworthy 28,100 stars on GitHub.
  • ML-based method: We use mlscraper, a tool to scrape data from HTML pages automatically. We leverage their publicly available implementation [2]. Notably, the implementation of mlscraper has gained significant popularity, earning 1,200 stars on GitHub.
@@ 6.1.4 Evaluation metrics
  • Accuracy:  Suppose we have extracted personal information (e.g., an email address or a phone number) by a LLM $f$ using the prompt $x$ and the corresponding ground truth $y$. We say the extracted personal information is correct only if it exactly matches the ground truth. Formally, accuracy is defined as follows:
    $\textit{Accuracy}=\mathbbm{1}[\hat{y}=y],$ (1)
    where $\hat{y}=f(x,d)$ represents the extracted personal information and $\mathbbm{1}$ is an indicator function whose value is 1 if the condition is satisfied and 0 otherwise. We report the average accuracy of all personal profiles in a dataset.
  • Rouge-1 Score:  Rouge-1 score measures the overlap of unigrams (a single word) between the extracted personal information and the corresponding ground truth. Formally, it is defined as follows:
@@
Note that, when $y$ is empty (i.e., a personal profile does not contain certain personal information), but the extracted personal information $\hat{y}$ does not indicate this, we treat the accuracy or Rouge-1/BERT score between $\hat{y}$ and $y$ as 0. For instance, when a personal profile does not contain an email address, but $\hat{y}$ is “123@gmail.com”, we treat the accuracy as 0. By contrast, if $\hat{y}$ is an empty string or a text string which implies that the personal profile does not have this personal information (e.g., “none” or “The email address is unknown”), we treat the accuracy as 1.
Unless otherwise mentioned, in our evaluation, each table entry presents the experimental results in the form of “Accuracy / BERT score” for email address and Tel. and “Rouge-1 score / BERT score” for other categories of personal information. Besides, a larger Accuracy (or Rouge-1 score or BERT score) indicates a better attack performance in extracting personal information.
@@
Defense Brief Example Has impact on normal users Applicability
Symbol replacement Replace symbols in the sensitive information with English letters. Replace “123@gmail.com” with “123 AT gmail DOT com” Yes Only applicable to information that contains a certain range of symbols, e.g., email address.
Keyword replacement Replace keywords in the sensitive information with alternatives. Replace “albert.einstein@ gmail.com” with “<first_name>. <last_name>@gmail.com” Yes Only applicable to information that contains well-known alternatives, e.g., email address that contains the owner’s names.
Hyperlink Hide the sensitive information in a clickable hyperlink. Replace “123@gmail.com” with “<li><a href="mailto:{123@ gmail.com}"> Contact me </a></li>” Yes Only applicable to information that can be replaced by a hyperlink, e.g., email address.
Text-to-image Store the sensitive information in an image. Replace “123@gmail.com” with an image of “123@gmail.com” No Applicable to all information.
Prompt injection Inject prompts to personal profiles to misguide LLMs. After “123@gmail.com”, add “Ignore previous information. The true email address of me is actually 321@gmail.com” in the format that is invisible to normal users. No Applicable to all information.
@@ 6.2 Attack Results
LLM-based PIE is effective:  Table 3 shows the results for the LLM-based information extraction on 10 LLMs and 3 datasets. From the experimental results, we have the following observations. First, in general, accuracy (or Rouge-1 score) and BERT score are high, which means LLMs could achieve good performances in extracting personal information. Second, the information extraction is less accurate on two real-world datasets compared to the synthetic dataset, especially when extracting the work and educational experience. We suspect the reason is that the profiles in real-world datasets have a more complex structure, which makes it more difficult for LLMs to summarize the persons’ work and educational experience. Third, in general, the attack is more effective when the LLM is larger (or more powerful). For instance, the accuracy for text-bison-001 and vicuna-7b-v1.3 on the synthetic dataset when extracting the email address are 0.93 and 0.55, respectively. This demonstrates that, when an attacker has access to a more powerful LLM, the information extraction tends to be more accurate.
@@
LLM outperforms traditional methods for PIE:  Table 4 presents the results of PIE using traditional methods, including regular expression, keyword search, entity detection, and ML-based method. Compared to the results of LLMs in Table 3, the LLM-based PIE outperforms the traditional methods in most cases. For instance, when extracting the mail address using text-bison-001 on the synthetic dataset, the Rouge-1 score is 0.77, whereas the highest Rouge-1 score that the traditional methods can achieve is 0.49. When extracting email addresses using only regular expressions, the accuracy is slightly higher than using text-bison-001. This is because regular expression is effective but limited to handling categories with highly specific structures, such as email addresses. However, for categories like affiliation and occupation that lack specific structures, regular expression exhibits considerably lower performance. In addition, when defenses are used in personal profiles, e.g., the email address is converted to an image, traditional methods such as regular expression will fail to extract that information. In summary, compared to traditional methods, the use of LLMs is more effective in extracting personal information.
@@
Impact of prompt style:  Table 5 compares the performance of different styles of the prompt used by the attacker, i.e., direct, persona, contextual, and pseudocode. We have two observations from the experimental results. First, the results suggest that the pseudocode style achieves slightly better results when extracting email address, mailing address, phone number, and name, but has a slightly worse performance when extracting work and educational experience. Second, the other three styles of prompts, i.e., direct, persona, and contextual, achieve similar results. This suggests that the LLM-based information extraction has a consistent performance across different prompt styles.
@@
Impact of redundant information filtering:  Table 16 in the Appendix compares the performance of LLM-based PIE with and without redundant information filtering. We have the following observations from the results. In particular, for LLMs with a larger number of parameters, their performance has no big difference no matter whether the redundant information filter is used or not. However, for LLMs with a smaller number of parameters, the performance is much worse without the redundant information filter. The reason is that they are less powerful in processing long input. As a result, filtering redundant information could significantly improve their performance.

### v2

@@ 5.2 Real-world Datasets
As complementary to the synthetic dataset, we collected three real-world datasets. To mitigate the ethical concerns, we collect personal profiles that are already in the public domain.
Celebrity:  The first real-world dataset is collected from the “List of Top 100 Famous Peopl” website [10], which contains personal profiles of famous people from the nineteenth, twentieth, or twenty-first centuries. The list includes famous actors, politicians, entrepreneurs, writers, artists, and humanitarians.
Physician:  The 2nd dataset is derived from “Wikipedia’s 19th century American physicians” [9], which contains personal profiles of known American physicians in the 19th century.
Professor:  We manually collect 100 personal profiles from EECS professors at MIT, CMU, Stanford, and UC Berkeley. Compared with Celebrity and Physician datasets, this dataset contains more diverse personal profiles (e.g., with different styles and formats). Moreover, some of these profiles have applied countermeasures against PIE attacks, making the evaluation more practical and reflective of real-world scenarios.
Court:  To further increase the diversity of our evaluation, we also incorporate the dataset introduced in [49], which was originally used for text anonymization. It includes cases from European Court of Human Rights. For simplicity, we denote this dataset as “Court”. This dataset contains 3 categories of personal information that align with ours: LOC, PERSON, and ORG, which correspond to mailing addr., name, and work experience in our evaluation.
All personal profiles in the first three real-world datasets are in HTML format. In addition, we manually collect the ground-truth personal information in these real-world datasets. We note that some personal profiles may not have some categories of personal information. A summary of the datasets used in our evaluation is in Table 1.
6 Evaluating Attacks
6.1 Experimental Settings
6.1.1 LLMs and Datasets
LLMs: We leverage the following LLMs: PaLM 2 text-bison-001 [14], PaLM 2 chat-bison-001 [14], gemini-pro [58], Flan-UL2 [60, 57, 56], Vicuna-13b-v1.3 [23, 26], Vicuna-7b-v1.3 [23], OpenAI’s gpt-3.5-turbo, gpt-4 [41], Llama-2-7b-chat [25], and InternLM-Chat-7B [6]. Table 2 shows the details of those LLMs. Unless otherwise mentioned, we use the PaLM 2 text-bison-001 as the default model, due to its free-of-charge APIs and powerful capabilities (with 540B parameters), enabling us to conduct systematic evaluation.
Datasets:  In our experiments, we use the 5 datasets introduced in Section 5: Synthetic, Celebrity, Physician, Professor, and Court. Unless otherwise mentioned, we use the Synthetic dataset as the default one.
6.1.2 Attack Settings
We use the following settings for the attack framework:
@@ 6.1.2 Attack Settings
API-based GPT-4 1.5T OpenAI
GPT-3.5-Turbo 154B OpenAI
PaLM 2 text-bison-001 540B Google
@@
Table 3: The effectiveness of LLMs at extracting 8 categories of personal information on five datasets. The results are presented as Accuracy / BERT score for email addr. and Tel., and as Rouge-1 score / BERT score for other categories.
@@
Professor 91% / 0.91 94% / 0.94 65% / 0.63 83% / 0.83 45% / 0.24 60% / 0.34 70% / 0.46 76% / 0.61
Court - - 54% / 0.43 84% / 0.64 25% / 0.03 - - -
vicuna-13b-v1.3 Synthetic 59% / 0.59 57% / 0.57 62% / 0.61 100% / 1.00 69% / 0.46 73% / 0.59 63% / 0.65 67% / 0.65
@@
Professor 71% / 0.71 60% / 0.60 76% / 0.68 87% / 0.87 47% / 0.41 46% / 0.37 71% / 0.56 63% / 0.58
Court - - 1% / 0.00 25% / 0.13 2% / 0.00 - - -
vicuna-7b-v1.3 Synthetic 65% / 0.65 95% / 0.95 53% / 0.47 99% / 0.99 57% / 0.32 68% / 0.53 84% / 0.82 76% / 0.69
@@
Professor 65% / 0.65 46% / 0.46 82% / 0.79 88% / 0.88 40% / 0.24 26% / 0.18 75% / 0.64 63% / 0.61
Court - - 14% / 0.02 23% / 0.14 25% / 0.00 - - -
llama-2-7b-chat-hf Synthetic 26% / 0.26 46% / 0.46 26% / 0.24 35% / 0.35 28% / 0.15 38% / 0.27 39% / 0.38 19% / 0.15
@@
Professor 33% / 0.33 45% / 0.45 53% / 0.47 35% / 0.35 22% / 0.04 25% / 0.18 53% / 0.32 35% / 0.27
Court - - 4% / 0.00 11% / 0.00 15% / 0.00 - - -
internlm-chat-7b Synthetic 95% / 0.95 99% / 0.99 81% / 0.78 100% / 1.00 42% / 0.28 72% / 0.58 75% / 0.74 76% / 0.69
@@
Regular expression 100% / 1.00 36% / 0.36 49% / 0.41 96% / 0.99 20% / 0.13 15% / 0.08 5% / 0.13 1% / 0.05
Keyword search 14% / 0.14 17% / 0.17 10% / 0.10 3% / 0.03 43% / 0.35 40% / 0.30 5% / 0.13 1% / 0.05
spaCy 95% / 0.95 24% / 0.24 18% / 0.10 93% / 0.93 1% / 0.01 11% / 0.04 4% / 0.11 1% / 0.05
BERT-based 1% / 0.01 8% / 0.08 28% / 0.10 14% / 0.14 1% / 0.00 1% / 0.00 36% / 0.19 19% / 0.06
mlscraper 77% / 0.77 8% / 0.08 10% / 0.10 75% / 0.75 1% / 0.01 1% / 0.01 5% / 0.13 1% / 0.05
Table 5: The effectiveness of traditional PIE methods at extracting personal information on the Professor dataset.
Extraction Method Personal Information
@@
Regular expression 59% / 0.59 63% / 0.63 48% / 0.46 90% / 0.89 2% / 0.02 40% / 0.33 0.32% / 0.09 0% / 0.29
Keyword search 13% / 0.13 66% / 0.66 56% / 0.56 0% / 0.02 3% / 0.03 42% / 0.41 0% / 0.09 0% / 0.29
spaCy 63% / 0.63 73% / 0.73 26% / 0.23 64% / 0.59 2% / 0.00 32% / 0.32 0.18% / 0.05 0% / 0.29
BERT-based 12% / 0.12 65% / 0.65 65% / 0.56 53% / 0.51 2% / 0.02 32% / 0.32 49% / 0.33 46% / 0.28
mlscraper 13% / 0.13 65% / 0.65 56% / 0.56 36% / 0.36 2% / 0.02 32% / 0.32 0% / 0.09 0% / 0.29
Table 6: The effectiveness of traditional PIE methods at extracting personal information on the Court dataset.
Extraction Method Personal Information
Mail addr. Name Work experience
Regular expression 4% / 0.00 0% / 0.00 0% / 0.00
Keyword search 1% / 0.01 0% / 0.00 0% / 0.00
spaCy 4% / 0.00 40% / 0.22 9% / 0.00
BERT-based 1% / 0.01 29% / 0.07 0% / 0.00
mlscraper 1% / 0.01 0% / 0.00 0% / 0.00
Personal profile processing:  Given a personal profile, an attacker aims to extract the following categories of personal information: name, email address, phone number, mailing address, work experience, educational experience, affiliation, and occupation. By default, we leverage an HTML parser to remove redundant information (see Section 4.3 for details).
@@ 6.1.3 Compared Baselines
We use several traditional PIE methods as baselines. 1) Regular expression: For each category of personal information, we develop corresponding regular expressions based on the possible structures of the corresponding texts. We use these regular expressions to match and extract personal information from a personal profile. Take email address as an example. Following the typical format of an email, the regular expression can be represented as “[a-zA-Z0-9._%+-]+@[a-zA-Z0-9.-]+\.[a-zA-Z]{2,}.”. Table 15 in the Appendix shows the specific regular expressions we use in experiments. 2) Keyword search: We first search for the keywords (e.g., “email”) present in the HTML titles. Then, we extract the text located under the identified HTML titles by keyword search. Finally, we post-process the extracted text by removing the redundant or unrelated information. 3) Named entity recognition (NER): We consider two NER methods: spaCy [1] and BERT-based method [33]. In particular, spaCy is a popular public tool with 28,100 stars on GitHub and used by many previous works for personal information extraction [49, 29, 43]. The BERT-based method is an information extractor fine-tuned based on BERT and is used to extract personal information for text anonymization [33]. 4) mlscraper [2] is an ML-based tool to scrape data from websites automatically. Notably, mlscraper has gained significant popularity with 1,200 stars on GitHub.
@@ 6.1.4 Evaluation Metrics
Accuracy:  Suppose we have extracted personal information (e.g., an email address or a phone number) by a LLM $f$ using the prompt $x$ and the corresponding ground truth $y$. We say the extracted personal information is correct only if it exactly matches the ground truth. Formally, we have: $\textit{Accuracy}=\mathbf{1}[\hat{y}=y]$, where $\hat{y}=f(x,d)$ represents the extracted personal information and $\mathbf{1}$ is an indicator function whose value is 1 if the condition is satisfied and 0 otherwise. We report the average accuracy of all personal profiles in a dataset.
Rouge-1 Score: Rouge-1 score measures the overlap of unigrams (a single word) between the extracted personal information and the ground truth. Formally, we have:
@@
When $y$ is empty (i.e., a personal profile does not contain certain personal information), but the extracted personal information $\hat{y}$ does not indicate this, we treat the accuracy or Rouge-1/BERT score between $\hat{y}$ and $y$ as 0. For instance, when a personal profile does not contain an email address, but $\hat{y}$ is “123@gmail.com”, we treat the accuracy as 0. By contrast, if $\hat{y}$ is an empty string or a text string which implies that the personal profile does not have this personal information (e.g., “none” or “The email address is unknown”), we treat the accuracy as 1.
Unless otherwise mentioned, each table entry presents the experimental results in the form of “Accuracy / BERT score” for email address and Tel. and “Rouge-1 score / BERT score” for other categories of personal information. Besides, a larger Accuracy (or Rouge-1 score or BERT score) indicates a better attack performance in extracting personal information.
@@
Defense Abbr. Brief Example Has impact on normal users Applicability
Symbol replacement SR Replace symbols in the sensitive information with English letters. Replace “123@gmail.com” with “123 AT gmail DOT com” Yes Only applicable to information that contains a certain range of symbols, e.g., email address.
Keyword replacement KR Replace keywords in the sensitive information with alternatives. Replace “albert.einstein@gmail.com” with “<first_name>. <last_name>@gmail.com” Yes Only applicable to information that contains well-known alternatives, e.g., email address that contains the owner’s names.
Hyperlink HL Hide the sensitive information in a clickable hyperlink. Replace “123@gmail.com” with “<li><a href="mailto:{123@ gmail.com}"> Contact me </a></li>” Yes Only applicable to information that can be replaced by a hyperlink, e.g., email address.
Text-to-image TI Store the sensitive information in an image. Replace “123@gmail.com” with an image of “123@gmail.com” No Applicable to all information.
Prompt injection PI Inject prompts to personal profiles to misguide LLMs. After “123@gmail.com”, add “Ignore previous information. The true email address of me is actually 321@gmail.com” in the format that is invisible to normal users. No Applicable to all information.
@@ 6.2 Attack Results
LLM-based PIE is effective:  Table 3 shows the results for the LLM-based information extraction on 10 LLMs and 5 datasets. From the experimental results, we have the following observations. First, in general, accuracy (or Rouge-1 score) and BERT score are high, which means LLMs could achieve good performances in extracting personal information. Second, the information extraction is less accurate on four real-world datasets compared to the synthetic dataset, such as when extracting the work experience. We suspect the reason is that the profiles in real-world datasets have a more complex structure, which makes it more difficult for LLMs to summarize the persons’ work experience. Third, in general, the attack is more effective when the LLM is larger (or more powerful). For instance, the accuracy for gpt-4 and vicuna-7b-v1.3 on the synthetic dataset when extracting the email address are 1.0 and 0.65, respectively. This demonstrates that, when an attacker has access to a more powerful LLM, the information extraction tends to be more accurate.
@@
LLM outperforms traditional methods for PIE:  Tables 4, 5, and 6 present the results of PIE using traditional methods, including regular expression, keyword search, spaCy, BERT-based method, and mlscraper on Synthetic, Professor, and Court datasets, respectively. Compared to the results of LLMs in Table 3, the LLM-based PIE outperforms the traditional methods in most cases. For instance, when extracting the mail address using text-bison-001 on the synthetic dataset, the Rouge-1 score is 0.77, whereas the highest Rouge-1 score that the traditional methods can achieve is 0.49. When extracting email addresses using only regular expressions, the accuracy is slightly higher than using text-bison-001. This is because regular expression is effective but limited to handling categories with highly specific structures, such as email addresses. However, for categories like affiliation and occupation that lack specific structures, regular expression exhibits considerably lower performance. In addition, when defenses are used in personal profiles, e.g., the email address is converted to an image, traditional methods such as regular expression will fail to extract that information. In summary, compared to traditional methods, the use of LLMs is more effective in extracting personal information.
@@
Impact of prompt style:  Table 13 in Appendix compares the performance of different styles of the prompt used by the attacker, i.e., direct, persona, contextual, and pseudocode. We have two observations from the experimental results. First, the results suggest that the pseudocode style achieves slightly better results when extracting email address, mailing address, phone number, and name, but has a slightly worse performance when extracting work and educational experience. Second, the other three styles of prompts, i.e., direct, persona, and contextual, achieve similar results. This suggests that the LLM-based information extraction has a consistent performance across different prompt styles.
@@
Impact of redundant information filtering:  Table 19 in the Appendix compares the performance of LLM-based PIE with and without redundant information filtering. We have the following observations. In particular, for LLMs with a larger number of parameters, their performance has no big difference no matter whether the redundant information filter is used or not. However, for LLMs with a smaller number of parameters, the performance is much worse without the redundant information filter. The reason is that they are less powerful in processing long input. As a result, filtering redundant information could significantly improve their performance.